Joe and Moe can mow the lawn in 26 minutes if they work together. If Moe works twice as fast as Joe, how long does it take Joe to mow the lawn alone?

Answer by nerdybill(7384)

Joe and Moe can mow the lawn in 26 minutes if they work together. If Moe works twice as fast as Joe, how long does it take Joe to mow the lawn alone?
.
Let J = time (minutes) Jo takes to mow alone
then
2j = time Moe takes
.
26(1/J + 1/(2J)) = 1
multiplying both sides by 2J:
26(2 + 1) = 2J
26(3) = 2J
[26(3)]/2 = J
[13(3)] = J
39 minutes = J (Jo's time)